OpenStudy (marigirl):

Integration application question: Please help me, Where can I start? I know it will be an exponential equation eventually. A farmer goes out at 7.30 am to check his stock and finds one of his cows dead in the creek. The temperature of the cow is 22 degrees Celsius and the temperature of the creek is 5 degrees Celsius. One hour later the temperature of the cow is 19 degrees Celsius. The normal body temperature of a healthy cow is 38.6 degrees Celsius. When did the cow die?

OpenStudy (anonymous):

Well, perhaps you can start with your model, the exponential decay function\[H(t) = T_0e^{-Rt}\]

OpenStudy (anonymous):

$$T_0$$ will probably be the cows temperature when the farmer found it dead and $$R$$ will be the rate at which heat dissipates from the cows body.

OpenStudy (marigirl):

t=0 cow's body temperature is 22 t=1 cow's body temperature is 19

OpenStudy (anonymous):

Yes. That allows you to compute R

OpenStudy (marigirl):

what is with the information about the creek being 5 degrees

OpenStudy (anonymous):

This is a good question, however, it seems to be erroneous information since the rate is fixed given the conditions at t = 0 and t = 1

OpenStudy (anonymous):

This is based on your model though.

OpenStudy (marigirl):

@freckles I would really appreciate your input :)

OpenStudy (marigirl):

im a bit lost with finding R @RBauer4

OpenStudy (anonymous):

Put t = 1, then we get\[22 e^{-R} = 19\]

ganeshie8 (ganeshie8):

morbid use of cows lol

OpenStudy (marigirl):

actually im very lost

OpenStudy (marigirl):

oh please help @ganeshie8 .. its bad enough that i have to read about deceased cows :(

OpenStudy (marigirl):

there is even a picture of a dead cow in the question!!!!! :( :( :( :( :( this is emotionally disturbing me

OpenStudy (anonymous):

The thing is, is that the farmer left it in the creek for another hour after he found it!

OpenStudy (marigirl):

and took its temperature

OpenStudy (anonymous):

Here is what I would do, take\[H(t) = 22 e^{-Rt}\]. Then the problem states that at t = 1, the cows temperature was 19, so equivalently \[H(1) = 22 e^{-R} = 19\]

OpenStudy (anonymous):

We can solve for R explicitly for \[e^{-R} = \frac{19}{22} \implies e^{R} = \frac{22}{19} \implies \ln(e^{R}) = R = \ln(22/19)\]

OpenStudy (anonymous):

Ok, now that we have R, all we need to do is put H(t) = 38.6 and find the value of t this corresponds to.

OpenStudy (freckles):

@marigirl are you ok with the equation @RBauer4 has left you with?

OpenStudy (marigirl):

it makes sense but i am still considering why the paddock temperature is stated. ill take a photo of the answer and send it to you guys

ganeshie8 (ganeshie8):

im getting 4:30 am using newton's law of cooling

OpenStudy (marigirl):

answer in the book states t=-3.51

ganeshie8 (ganeshie8):

right, they are using newton's law of cooling do you want to know how to setup the differential equation and solve it

OpenStudy (marigirl):

yes please i will also upload the model answer shown

OpenStudy (marigirl):

ANSWER

ganeshie8 (ganeshie8):

Let \(y(t)\) represent the temperature of cow \(t\) hours after \(7:30\) am, (time after it was observed for the first time) then the temperature in cow follows the newton's law of cooling: \[y' = k(5-y)\tag{1}\]

ganeshie8 (ganeshie8):

does that look familiar to you if not, you may replace \(y\) by \(T\)

OpenStudy (marigirl):

thanks, i am not seeing it now, but i will think about it.

ganeshie8 (ganeshie8):

familiar with separation of variables to solve differential equation ?

OpenStudy (anonymous):

Not that I have never solved such an ODE before, but I was never taught that this particular ODE models cooling. Thanks for the clarification @ganeshie8

ganeshie8 (ganeshie8):

question should explicitly specify the model to use i guess because there are several models...